Please adhere to the rounding instructions(in red) for calculations and final answer. If the answer is not exact, it will be incorrect. Thanks!

Question 11 3 pts Your firm has been working on an advanced technology. This technology will be available in the near term. The firm anticipates the first annual cash flow from the technology to be $141504, received three years from today. Subsequent annual cash flows will grow at 2.33% in perpetuity. What is the present value of the technology if the discount rate is 8.35%? (Round answer to 2 decimal places. Do not round intermediate calculations) Topic: Discounted Cash Flow Valuation
Answer #1

PV= sum of present value od all future cash flows

CF1= 141504

CF2= 141504(1+g)^1= CF1(1.0233)

CF3 =141504(1+g)^2= CF1(1.0233)^2

this goes on till perpetuity(Infinity)

Now discounting these to T=0

PV= CF1/(1+r)^3 + CF2/(1+r)^4 +CF3/(1+r)^5 ....... Infinity --- This is an infinte GP with first term(t1) as CF1/(1+r)^3 and ratio as (1+g)/(1+r)

PV= t1/(r-g)

t1= 141504/(1.835)^3= 11057.8125

PV= 11057.8125/(0.0835-0.0233)= 183684.59 ---- ANSWER